OpenStudy (anonymous):

Which rate is the best deal for a package of peanuts? A. $2.05 for 75 g B. $2.25 for 80 g C. $1.40 for 50 g D. $1.02 for 35 g

OpenStudy (anonymous):

fans and medals

OpenStudy (anonymous):

calculate the rate for per gram = weight/cost the one with the lowest rate is the best deal

OpenStudy (anonymous):

so which is it?

OpenStudy (anonymous):

@sangya21?

OpenStudy (anonymous):

which one do you think?

OpenStudy (anonymous):

b

OpenStudy (anonymous):

Whats the rate for each of the peanut

OpenStudy (anonymous):

O.O

OpenStudy (anonymous):

:-S

OpenStudy (anonymous):

he/she told you exactly what to do, divide weight by cost and the smallest number is the best deal

OpenStudy (anonymous):

1. rate = 2.05/75 2. rate = 2.25/80 3. rate = 1.40/ 50 4. rate = 1.02/35 sorry it should be cost per gram

OpenStudy (anonymous):

oh, okay but who do i give the medal to

OpenStudy (anonymous):

whomever you want to. main purpose was solving the question such that you can get the idea of it

OpenStudy (anonymous):

him

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Dear @Audeezy, its her.

OpenStudy (anonymous):

oo, sorry